Rank these species by their ability to act as an oxidizing agent. Cl2, I2, Mg2+, Fe2+
Answer #1

first consider the standard reduction potentials

Eo Mg+2/Mg = -2.372

Eo Fe+2 / Fe = -0.44

Eo Cl2 / 2Cl- = 1.36

Eo I2/2I- = 0.54

to act as oxidizing agent , the reagent must undergo reduction

so

higher the reduction potential , better the oxidizing agent

so

the order is

Cl2 > I2 > Fe+2 > Mg+2